# GPU-accelerated LISA parameter estimation with full time domain response

### v2

@@ II.1 Waveform generation
Results indicate that the new Python implementation of IMRPhenomT provides significant computational advantages. Specifically, it outperforms the original C-based LAL implementation and achieves performance comparable to the Fourier-domain models on the CPU. In its GPU version, the Python implementation demonstrates a two-order-of-magnitude improvement over the LAL IMRPhenomT and a one-order-of-magnitude improvement over IMRPhenomX, the latter being the most computationally efficient waveform family available thus far. It is also worth noting that our Python implementation of IMRPhenomT does not yet incorporate any non-uniform grid acceleration techniques, such as those used in IMRPhenomX with the Multibanding technique. Porting this technique to the time domain could reduce even more the computational cost.
Given the evaluation speed-up factor introduced with the new implementation and its enhanced simplicity compared to LALSuite, there are efforts to also port the IMRPhenomX family and incorporate new developments like the (2,0) mode with memory Rosselló-Sastre et al. (2024) and eccentricity effects in the time domain de Lluc Planas (2024) and the Fourier domain Ramos-Buades (2024).
@@
The considerable speed-up and simplified structure provided by this new Python implementation motivates additional efforts to port the entire IMRPhenomX waveform family into Python as well. Future work includes implementing new features such as the (2,0) mode with memory effects Rosselló-Sastre et al. (2024) and eccentricity effects in the time domain de Lluc Planas (2024) as well as in the Fourier domain Ramos-Buades (2024). These enhancements aim to expand the versatility and precision of IMRPhenom models in gravitational wave analysis across a wider array of astrophysical scenarios, benefiting applications in both LISA and ground-based detectors.
@@ II.2 Time domain LISA response
For current ground-based detectors, the response to incoming signals is computed simply multiplying the two polarizations by the well-known antenna patterns, which can be treated as static due to the short duration of the signals. However, in LISA, signals can span over a much longer time, necessitating a dynamical time-varying detector response. More importantly, the response from each link needs to be appropriately combined through TDI to suppress the laser frequency noise by several orders of magnitude. Without TDI, detecting gravitational wave signals under such significant noise would be nearly impossible. Working directly in the time domain allows for straightforward testing of various TDI configurations and the incorporation of realistic orbits with unequal arm lengths, capabilities that are currently unattainable with the Fourier domain formalism. Here we outline the method to compute LISA’s time-dependent response, closely following the approach and notation established in the LISA Data Challenge (LDC) LISA Data Challenge working group (2022); S. Babak, A. Petiteau (2020); S. Babak, M. Le Jeune, A. Petiteau, M. Vallisneri (2020).
@@
To compute the LISA response, the positions of the sender and receiver ($\textbf{r}_{s}$, $\textbf{r}_{r}$) the light travel time in each link ($L_{l}$) must be tracked over the whole observation period ($t_{\mathrm{obs}}$). These quantities can be obtained from multiple simulation codes of the LISA orbital dynamics and only need to be computed once. We will compare two types of orbits, one with equal arms computed with the AnalyticOrbit class provided in the LISA Data Challenge tools LISA Data Challenge working group (2022) and a second one with the realistic orbits provided by ESA Martens and Joffre (2021).
@@
k $=\big{(}-\cos(\beta)\cos(\lambda),-\cos(\beta)\sin(\lambda),-\sin(\beta)\big{)}$ (7)
v $=\big{(}-\sin(\beta)\cos(\lambda),-\sin(\beta)\sin(\lambda),\cos(\beta)\big{)}$
u $=\big{(}\sin(\lambda),-\cos(\lambda),0\big{)},$
@@
The other two Michelson variables $Y,Z$ are obtained by simple and double cyclic permutations of the $slr$ indices respectively. Independent channels can be constructed under the premise of identical and uncorrelated noise across the detector arms as
@@
$\begin{split}\log\mathcal{L}=-\big{(}&(\tilde{A}_{inj}|\tilde{A})-\frac{1}{2}(\tilde{A}|\tilde{A})\\ +&(\tilde{E}_{inj}|\tilde{E})-\frac{1}{2}(\tilde{E}|\tilde{E})\\ +&(\tilde{T}_{inj}|\tilde{T})-\frac{1}{2}(\tilde{T}|\tilde{T})\big{)}.\end{split}$ (12)
@@
[…]
@@
$S_{\mathrm{acc}}$ $=\left(2.4\cdot 10^{-12}\right)^{2}\>\left(1+\left(\frac{4\cdot 10^{-4}}{f}\right)^{2}\right)\left(1+\left(\frac{f}{0.008}\right)^{4}\right)\cdot$ (14)
@@
[…]
@@ II.3 Injections setup
In the most general scenario, one would use the full LISA frequency range $0.110^{-4}-$ 0.1 Hz, which translates to a sampling time interval of $\Delta t=1/(2f_{max})=5s$. Here we will generate the waveforms with a minimum frequency of $10^{-4}Hz$, but sample instead every $\Delta t=10s$ and then linearly interpolate them to generate the arm and TDI responses with a sampling time of $5s$ for a faster evaluation. Once the Fourier transform is performed, we will restrict the frequency range to $0.023\cdot 10^{-4}-$ 0.02 Hz to expedite the inner products calculation. The injected signal will represent an MBHB with a duration of approximately two months and will be merged well below the maximum frequency of 0.02 Hz. We will inject this system with and without precessing spins under zero-noise [^4] The time series representing the injected signal, designed to mimic the detector’s data, comprises solely the gravitational wave template, with no noise time series added. conditions in the signal. The parameters to be injected, along with the corresponding uniform prior ranges, are detailed in Tab. 1.
@@
For the injection and recovery processes, we will utilize the same waveform model and same TDI configurations. Bayesian inference will be executed using the Bilby package Ashton et al. (2019), which is extensively used for analyzing data from LVK collaboration. Notably, no modifications to the Bilby package are necessary to interface with our new waveform generation and LISA response calculations, as this functionality has been integrated into the waveform generator. A preliminary profiling test for the aligned-spin case revealed that the Fourier transform, arm response, and TDI calculation contribute approximately to the 10% of the total computational cost. In contrast, waveform model generation accounts for 20-40%, while likelihood calculation represents 50-70%. This substantial computational cost associated with the likelihood calculation has been identified, and a solution will be provided in a future update of our code. The goal is to ensure that the primary computational cost is attributed to the waveform model generation.
@@
We conducted preliminary test runs in which a reduced number of parameters were sampled, and these runs successfully recovered all the parameters with Gaussian-shaped posteriors. However, as we increased the number of parameters in the precessing case, the three main samplers we tested (ptemcee Vousden et al. (2015), dynesty Speagle (2020), and multinest Feroz et al. (2009)) did not converge within a time frame of a few days, due to the high computational cost and higher dimensionality of the parameter space. In order to speed up the results, the injections presented in the next section will not commence with a random sampling across the parameter space but will start from a region near the injected values based on Fisher estimates. This Fisher initialization method generates random starting points drawn from a multivariate Gaussian distribution where the mean is given by the injected values and the covariance matrix is the inverse of the Fisher matrix. This initialization method is also used in other pipelines like lisabeta Marsat et al. (2021). The Fisher matrix is computed from the derivatives of the likelihood function, using first-order finite differences as implemented in the calculate_FIM function within Bilby.
@@
The number of initial points for the sampler corresponds to the number of live points for the dynesty sampler and the product of the number of walkers and temperatures for ptemcee. Despite this initialization, we found that the dynesty sampler struggled more than ptemcee to achieve convergence, prompting us to focus exclusively on the results obtained with ptemcee. For our analysis, we utilized 28 walkers and 4 temperatures for all the runs, executing the computations on a single Nvidia Tesla V100 GPU.
@@
Another approach to enhance the convergence speed of the sampler, is to employ Fisher priors, which utilize the diagonal values of the covariance matrix as standard deviations for univariate Gaussian distributions assigned as priors for each parameter. Both methods necessitate prior knowledge of the true parameters or at least values in close proximity to them. This scenario is analogous to analyses utilizing heterodyne likelihood Cornish (2021) or relative binning Zackay et al. (2018); Leslie et al. (2021); Krishna et al. (2023) algorithms, both of which require a fiducial point situated near the maximum likelihood estimate. Such a point can often be derived from the best-fitting template obtained via a matched-filter search, where the injected signal is compared against a pre-generated bank of gravitational wave templates. In-depth investigations are required to assess how effectively these estimates perform across the parameter space, the necessary proximity to the true values, and whether estimates derived from dominant-mode, non-precessing template banks can be applied to analyses involving subdominant harmonics and precession.
@@
In the following section, we demonstrate that using Fisher initialization allows us to accurately recover all injected parameters within a few hours. To validate our estimates, we will compare our Bayesian posteriors with those derived from a full Fisher analysis. The Fisher analysis serves as a useful approximation for parameter estimation, achieving exact results in the limit of infinite SNR Cutler and Flanagan (1994); Cutler and Vallisneri (2007); Vallisneri (2008). Consequently, as the SNR increases, we would expect our Bayesian inference posteriors to approach the Fisher results. At finite SNR, our posteriors should be less informative, appearing “wider” compared to those from the Fisher analysis. This relationship will be examined in Sec. III.2. The Fisher posteriors are obtained by drawing $10^{4}$ samples from the gaussian multivariate distribution previously described. In future work, we aim to investigate the potential for running without sampler initialization by further optimizing the likelihood evaluation process. For now, the scenarios presented here serve as a critical testbed, providing valuable insights into LISA’s capabilities for precise parameter measurements across various scenarios.
@@
minimum frequency [Hz] $10^{-4}$
reference frequency [Hz] $10^{-4}$
@@ III Results
We present a set of zero-noise injections to assess LISA’s capacity to measure source parameters for both aligned-spin and precessing systems. We examine the effect of including subdominant harmonics on parameter recovery, compare the Bayesian posterior distributions against the approximate Fisher analyses, and explore the influence of unequal arms, different TDI channels and TDI generations on the results. These comparisons are among the first to utilize full Bayesian inference, providing valuable insights into the benefits of incorporating additional physical effects in waveform models and understanding how various LISA configurations may impact the accuracy of parameter recovery.
Bayesian inference is carried out employing the Bilby package in serial mode, with the ptemcee sampler initialized via Fisher matrices to expedite convergence. All the runs employ equal arm configuration except those in Sec. III.3.
In this initial study, both the signal injection and parameter recovery are performed with exactly the same waveform model and LISA configuration. A more realistic approach, however, would involve using a fixed injected signal (such as an NR injection) and then conducting parameter recovery with various waveform models (e.g., IMRPhenomT, IMRPhenomTHM, IMRPhenomTPHM…) and different LISA settings. However, due to the higher computational demand of this setup, such an analysis is deferred to future work.
@@
The median values, the 90% credible intervals, and relative errors with respect to the injected values [^5] Relative errors are computed dividing the width of the 90% credible intervals by the injected value. are summarized in Tabs. 2 and 3, alongside the total number of samples and sampling times for each run.
@@ III.1 Effect of subdominant harmonics
Incorporating additional physical effects in gravitational waveforms generally enhances the accuracy of source parameter estimation by breaking degeneracies between parameters. In particular, subdominant harmonics are crucial to break the degeneracy between inclination and distance London et al. (2018) and, to a lesser extent, between the polarization angle and reference phase. Distance impacts the waveform by overally scaling its amplitude. For the dominant mode, the inclination angle entering in the spin-2 weighted spherical harmonics ($Y_{22}$, $Y_{2-2}$) is also an overall scaling factor, therefore degenerated with the distance. With subdominant harmonics included, the inclination affects each individual harmonic, changing their relative contribution to the multimode waveform, thus breaking the degeneracy.
Similarly, the polarization angle $\psi$ is an overall phase offset (multiplying the waveform $h_{+}-ih_{\times}$ by $e^{i2\psi}$). For a waveform containing only the dominant mode, this resembles the influence of the reference phase, as it enters through the complex exponential of the $Y_{lm}$ as $e^{im\phi}$. With subdominant harmonics, the reference phase affects differently each harmonic, thus breaking the degeneracy similarly to the previous case.
These two degeneracy breakings can be observed very prominently in Fig. 2 for the aligned-spin case. The measurements of distance, inclination, reference phase and polarization are much more stringent with subdominant harmonics. This observation aligns with the findings reported in Marsat et al. (2021) also for LISA parameter estimation.
The accuracy of the rest of the parameters also improves including subdominant modes. In Tab. 2 we show all the recovered parameters with their 90% confidence intervals and relative errors with respect to the injected values. Notably, unlike current ground-based detectors, LISA demonstrates the ability to accurately measure individual spin components, which is done at the 1% level. This capability is crucial for population analysis and aids to discern between different formation channels. Furthermore, we emphasize the importance of incorporating subdominant harmonics for achieving a more precise determination of the source’s sky location, which is essential for multimessenger follow-up observations.
@@
When comparing results with and without precession, we observe no substantial improvement in the accuracy of recovered parameters (see Tab. 2). Initially, we hypothesized that this limited impact might stem from the injected system’s orientation, as it was near to face-off ($\theta_{JN}=153^{\circ}$); thus, we conducted additional analyses on an edge-on system but again observed no significant improvement. In fact, if we compare the posteriors from the Fisher analysis in the next section, we see that parameters like masses, distance and sky location have a noticeable accuracy improvement when including precession. This could suggest a need for enhanced sampling configurations to adequately navigate the higher-dimensional parameter space introduced by precessin systems. In this study, the sampler settings were held constant across all runs, without in-depth tuning for precessing cases, which may limit parameter space exploration. Furthermore, as we are comparing recoveries of different injected signals (a non-precessing one versus a precessing one) our initial expectation that precession would consistently enhance parameter recovery may not fully hold in this context. A comprehensive, systematic investigation is planned for future work. Here, our main aim is to showcase illustrative examples of potential outcomes with the newly developed framework.

### v3

@@ II.1 Waveform generation
Results indicate that the new Python implementation of IMRPhenomT provides significant computational advantages. Specifically, it outperforms the original C-based LAL implementation and achieves performance comparable to the Fourier-domain models on the CPU. In its GPU version, the Python implementation demonstrates a two-order-of-magnitude improvement over the LAL IMRPhenomT and a one-order-of-magnitude improvement over IMRPhenomX, the latter being the most computationally efficient waveform family available so far. It is also worth noting that our Python implementation of IMRPhenomT does not yet incorporate any nonuniform grid acceleration techniques, such as those used in IMRPhenomX with the Multibanding technique. Porting this technique to the time domain could reduce the computational cost even more.
@@
The considerable speed-up and simplified structure provided by this new Python implementation motivates additional efforts to port the entire IMRPhenomX waveform family into Python as well. Ongoing work includes implementing new features such as the (2,0) mode with memory effects Rosselló-Sastre et al. (2024) and eccentricity effects in the time domain de Lluc Planas et al. (2025) as well as in the Fourier domain Ramos-Buades (2025). These enhancements aim to expand the versatility and precision of IMRPhenom models in gravitational wave analysis across a wider array of astrophysical scenarios, benefiting applications in both LISA and ground-based detectors.
@@ II.2 Time domain LISA response
For current ground-based detectors, the response to incoming signals is computed simply multiplying the two polarizations by the well-known antenna patterns, which can be treated as static due to the short duration of the signals. However, in LISA, signals can span over a much longer time, necessitating a dynamical time-varying detector response. More importantly, the response from each link needs to be appropriately combined through TDI to suppress the laser frequency noise by several orders of magnitude. Without TDI, detecting gravitational wave signals under such significant noise would be nearly impossible. Working directly in the time domain allows for straightforward testing of various TDI configurations and the incorporation of realistic orbits with unequal-arm lengths, capabilities that are currently unattainable with the Fourier domain formalism. Here we outline the method to compute LISA’s time-dependent response, closely following the approach and notation established in the LISA Data Challenges (LDC) LISA Data Challenge working group (2022); S. Babak, A. Petiteau (2020); S. Babak, M. Le Jeune, A. Petiteau, M. Vallisneri (2020).
@@
To compute the LISA response, the positions of the sender and receiver ($\textbf{r}_{s}$, $\textbf{r}_{r}$) and the light travel time in each link ($L_{l}$) must be tracked over the whole observation period ($t_{\mathrm{obs}}$). These quantities can be obtained from multiple simulation codes of the LISA orbital dynamics and only need to be computed once. In Sec. III.3 we will compare two types of orbits, one with equal-arms computed with the AnalyticOrbit class provided in the LISA Data Challenge tools LISA Data Challenge working group (2022) and a second one with realistic unequal-arm orbits provided by ESA Martens and Joffre (2021).
@@
k $=\big{(}-\cos\beta\cos\lambda,-\cos\beta\sin\lambda,-\sin\beta\big{)}$ (7)
v $=\big{(}-\sin\beta\cos\lambda,-\sin\beta\sin\lambda,\cos\beta\big{)}$
u $=\big{(}\sin\lambda,-\cos\lambda,0\big{)},$
@@
The other two Michelson variables $Y$, $Z$ are obtained by simple and double cyclic permutations of the slrslr indices respectively. Independent channels can be constructed under the premise of identical and uncorrelated noise across the detector arms as
@@
[…]
@@
$(\tilde{a}|\tilde{b})=4\mathrm{Re}\int_{0}^{\infty}\frac{\tilde{a}(f)\tilde{b}^{*}(f)}{S_{n}(f)}df.$ (13)
@@
$S_{\mathrm{acc}}$ $=\left(2.4\cdot 10^{-12}\right)^{2}\>\left(1+\left(\frac{4\cdot 10^{-4}}{f}\right)^{2}\right)\left(1+\left(\frac{f}{0.008}\right)^{4}\right)\times$ (14)
@@
$S_{\mathrm{op}}$ $=\left(7.9\>10^{-12}\right)^{2}\left(1+\left(\frac{0.002}{f}\right)^{4}\right)\left(\frac{2\pi f}{c}\right)^{2}.$
@@ II.3 Injections setup
In the most general scenario, one would use the full LISA frequency range $10^{-4}-0.1$ Hz, which translates to a sampling time interval of $\Delta t=1/(2f_{max})=5s$. Here we will generate the waveforms with a minimum frequency of $10^{-4}$ Hz, but sample instead every $\Delta t=10s$ and then linearly interpolate them to generate the arm and TDI responses with a sampling time of $5$ s for a faster evaluation. Once the Fourier transform is performed, we will restrict the frequency range to $3\cdot 10^{-4}-0.02$ Hz to expedite the inner products calculation. The injected signal will represent an MBHB with a duration of approximately two months and will be merged well below the maximum frequency of 0.02 Hz. We will inject this system with and without precessing spins under zero-noise [^4] The time series representing the injected signal, designed to mimic the detector’s data, comprises solely the gravitational wave template, with no noise time series added. conditions in the signal. The parameters to be injected, along with the corresponding uniform prior ranges, are detailed in Tab. 1.
@@
For the injection and recovery processes, we will utilize the same waveform model and same TDI configurations. Bayesian inference will be executed using the Bilby package Ashton et al. (2019), which is extensively used for analyzing data in the LVK collaboration. Notably, no modifications to the Bilby package are necessary to interface with our new waveform generation and LISA response calculations, as this functionality has been integrated into the waveform generator. A preliminary profiling test for the aligned-spin case revealed that the Fourier transform, arm response, and TDI calculation contribute approximately to the 10% of the total computational cost. In contrast, waveform model generation accounts for 20-40%, while likelihood calculation represents 50-70%. This substantial computational cost associated with the likelihood calculation has been identified, and a solution will be provided in a future update of our code. The goal is to ensure that the primary computational cost is attributed solely to the waveform model generation.
@@
We conducted preliminary test runs in which a reduced number of parameters were sampled, and these runs successfully recovered all the parameters with Gaussian-shaped posteriors. However, as we increased the number of parameters in the precessing case, the three main samplers we tested (ptemcee Vousden et al. (2015), dynesty Speagle (2020), and multinest Feroz et al. (2009)) did not converge within a time frame of a few days, due to the high computational cost and the high dimensionality of the parameter space. In order to speed up the results, the injections presented in the next section will not commence with a random sampling across the parameter space but will start from a region near the injected values based on Fisher estimates. This Fisher initialization method generates random starting points drawn from a multivariate Gaussian distribution where the mean is given by the injected values and the covariance matrix by the inverse of the Fisher matrix. This initialization method is also used in other pipelines like lisabeta Marsat et al. (2021). The Fisher matrix is computed from the derivatives of the likelihood function, using first-order finite differences as implemented in the calculate_FIM function within Bilby.
@@
The number of initial points provided to the sampler corresponds to the number of live points for dynesty and the product of the number of walkers and temperatures for ptemcee. Despite this initialization, we found that the dynesty sampler struggled more than ptemcee to achieve convergence, prompting us to focus exclusively on the results obtained with ptemcee. For our analysis, we utilized 28 walkers and 4 temperatures for all the runs, executing the computations on a single Nvidia Tesla V100 GPU.
@@
Another approach to enhance the convergence speed of the sampler is to employ Fisher priors, which utilize the diagonal values of the covariance matrix as standard deviations for univariate Gaussian distributions assigned as priors for each parameter. Both methods necessitate prior knowledge of the true parameters or at least values in close proximity to them. This scenario is analogous to analyses utilizing heterodyne likelihood Cornish (2021) or relative binning Zackay et al. (2018); Leslie et al. (2021); Krishna et al. (2023) algorithms, both of which require a fiducial point situated near the maximum likelihood estimate. Such a point can often be derived from the best-fitting template obtained via a matched-filter search, where the injected signal is compared against a pre-generated bank of gravitational wave templates. In-depth investigations are required to assess how effectively these estimates perform across the parameter space, the necessary proximity to the true values, and whether estimates derived from dominant-mode, non-precessing template banks can be applied to analyses involving subdominant harmonics and precession.
@@
In the following section, we demonstrate that using Fisher initialization allows us to accurately recover all injected parameters within a few hours. To validate our estimates, we will compare our Bayesian posteriors with those derived from a full Fisher analysis. The Fisher analysis serves as a useful approximation for parameter estimation, achieving exact results in the limit of infinite SNR Cutler and Flanagan (1994); Cutler and Vallisneri (2007); Vallisneri (2008). Consequently, as the SNR increases, we would expect our Bayesian inference posteriors to approach the Fisher results. At finite SNR, our posteriors should be less informative, appearing “wider” compared to those from the Fisher analysis. This relationship will be examined in Sec. III.2. The Fisher posteriors are obtained by drawing $10^{4}$ samples from the Gaussian multivariate distribution previously described. In future work, we will investigate the potential for running without sampler initialization by further optimizing the likelihood evaluation process. For now, the configuration presented here serves as a critical testbed, providing valuable insights into LISA’s capabilities for precise parameter measurements across various scenarios.
@@
minimum frequency [Hz] $10^{-4}$ -
reference frequency [Hz] $10^{-4}$ -
@@ III Results
We present a set of zero-noise injections to assess the LISA’s capacity to measure source parameters for both aligned-spin and precessing systems. We examine the effect of including subdominant harmonics on parameter recovery, compare the Bayesian posterior distributions against the approximate Fisher analyses, and explore the influence of unequal-arms, different TDI channels and TDI generations on the results. These comparisons are among the first to utilize full Bayesian inference, providing valuable insights into the benefits of incorporating additional physical effects in waveform models and understanding how various LISA configurations may impact the accuracy of parameter recovery.
Bayesian inference is carried out employing the Bilby package in serial mode, with the ptemcee sampler initialized via Fisher matrices to expedite convergence. All the runs employ an equal-arm configuration except those in Sec. III.3.
In this initial study, both the signal injection and parameter recovery are performed with exactly the same waveform model and LISA configuration. A more realistic approach would use a fixed injected signal (such as an NR simulation) and then conduct parameter recovery with various waveform models (e.g., IMRPhenomT, IMRPhenomTHM, IMRPhenomTPHM…) and different LISA settings. However, due to the higher computational demand of this setup, such analysis is deferred to future work.
@@
The median values, the 90% credible intervals, and the relative errors with respect to the injected values [^5] Relative errors are computed dividing the width of the 90% credible intervals by the injected value. are summarized in Tabs. 2 and 3, alongside the total number of samples and sampling times for each run.
@@ III.1 Effect of subdominant harmonics
Incorporating additional physical effects in gravitational waveforms generally enhances the accuracy of source parameter estimation by breaking degeneracies between parameters. In particular, subdominant harmonics are crucial to break the degeneracy between inclination and distance London et al. (2018) and between the polarization angle and the reference phase.
Distance impacts the waveform by overally scaling its amplitude. For the dominant mode, the inclination angle entering in the spin-2 weighted spherical harmonics ($Y_{22}$, $Y_{2-2}$) is also an overall scaling factor, therefore degenerated with the distance. With subdominant harmonics included, the inclination affects each individual harmonic, changing their relative contribution to the multimode waveform, thus breaking the degeneracy. Similarly, the polarization angle $\psi$ is an overall phase offset (multiplying the waveform $h_{+}-ih_{\times}$ by $e^{i2\psi}$). For a waveform containing only the dominant mode, this resembles the influence of the reference phase, as it enters through the complex exponential of the $Y_{lm}$ as $e^{im\phi}$ and $m=2$. With subdominant harmonics, the reference phase affects each harmonic differently, thus breaking the degeneracy similarly to the previous case.
These two degeneracy breakings can be observed very prominently in Fig. 3 for the aligned-spin case. The measurements of distance, inclination, reference phase and polarization angle are much more stringent with subdominant harmonics. This observation aligns with the findings reported in Marsat et al. (2021) also in the context of LISA parameter estimation.
The accuracy of the rest of the parameters also improves including subdominant modes. In Tab. 2 we show all the recovered parameters with their 90% confidence intervals and relative errors with respect to the injected values. Notably, unlike current ground-based detectors, LISA demonstrates the ability to accurately measure individual spin components, which is achieved at the 1% level. This capability is crucial for population analysis and aids in discerning between different formation channels. Furthermore, we emphasize the importance of incorporating subdominant harmonics to achieve a more precise determination of the source’s sky location, which is essential for multimessenger follow-up observations.
@@
When comparing results with and without precession, we observe no substantial improvement in the accuracy of recovered parameters (see Tab. 2). Initially, we hypothesized that this limited impact might stem from the injected system’s orientation, as it was close to face-off ($\theta_{JN}=153^{\circ}$); thus, we conducted additional analyses on an edge-on system but again observed no significant improvement. In fact, if we compare the posteriors from the Fisher analysis in the next section, we see that parameters like masses, distance and sky location have a noticeable accuracy improvement when including precession. This could suggest a need for enhanced sampling configurations to adequately navigate the higher-dimensional parameter space introduced by precessing systems. In this study, the sampler settings were held the same across all the runs, without in-depth tuning for precessing cases, which may limit parameter space exploration. Furthermore, as we are comparing recoveries of different injected signals (a non-precessing one versus a precessing one) our initial expectation that precession would consistently enhance parameter recovery may not fully hold in this context. A comprehensive, systematic investigation is planned for future work. Here, our main aim is to showcase illustrative examples of potential outcomes with the newly developed framework.